OpenStudy (anonymous):

4x+3y=55 an 5x-3y=35 substitution method???

OpenStudy (anonymous):

-3y=35-5x 3y=5x-35 y=(5x-35)/3 substitute in 1st eqn 4x+3*((5x-35)/3)=55 4x+5x-35=55 9x=90 x=10 substitute x=10 40+3y=55 3y=15 y=5